OpenStudy (anonymous):

find the graph for the inequality y > -4x - 8

jimthompson5910 (jim_thompson5910):

are you able to graph y = -4x - 8 ?

OpenStudy (anonymous):

I'm not sure, to be honest. I haven't taken math in a long time, so I don't really remember the formulas to do any of this.

jimthompson5910 (jim_thompson5910):

if x = 0, then what is y equal to? use y = -4x - 8

jimthompson5910 (jim_thompson5910):

if you aren't sure, I'll show you y = -4x - 8 y = -4(0) - 8 ... replace x with 0 y = 0 - 8 y = -8 so if x = 0, then y = -8 we now have the ordered pair (0, -8)

jimthompson5910 (jim_thompson5910):

hopefully that makes sense?